OpenStudy (anonymous):

What is the sum of a 34-term arithmetic sequence where the first term is 14 and the last term is 146? 2,550 2,720 2,890 2,982

OpenStudy (lgbasallote):

use the equation \[\large S_n = \frac{n(A_1 + A_n)}{2}\]

OpenStudy (lgbasallote):

where Sn is the sum of the series n = number of terms in the series A1 = first term An = last term

OpenStudy (anonymous):

2,720?

OpenStudy (anonymous):

Okay, this one's a little more abstract. We have the common difference is d = (a34 - a1)/33 = (146 - 14)/33 = 4 Now a1 = a1 a2 = a1 + d a3 = a1 + 2d ... a34 = a1 + 33d So to add all these up we need to add the first 33 natural numbers together. the formula is 1 + 2 + ... + n = n(n+1)/2 Which I'll explain if you'd like. Using this formula we have that the sum from a1 to a34 is 34*a1 + (33(34)/2)d = 34*14 + 33*17*4 = 476 + 2244 = 2720

OpenStudy (anonymous):

Okay, well, there's a longwinded explanation of how it works

jimthompson5910 (jim_thompson5910):

yes you have it right kerber123